OpenStudy (anonymous):

positive charge Q is distributed uniformly throughout an insulting sphere of radius R, centered at the origin. A particle with positive charge Q is place at x = 2R on the x axis. what is the magnitude of the electric field at x = R/2 on the x axis?

OpenStudy (anonymous):

when ever they say uniform charge distribution, there can be no electric field inside the material. Now use guass's law to find the electric field which can be showed to be \[q/(4*\pi*\epsilon_0*R)\] Substitute R= R/2 in this expression to get the answer.